Data Table 1 (Click on the icon located on the top-right corner of the data table below in order to copy its contents into a spreadsheet) Alternative Single amount $27,500 at the end of 3 years $53,500 at the end of 9 years $172,000 at the end of 20 years PrintDoneTime value comparisons of single amounts Personal Finance Problem In exchange for a $23,000 payment today, a well-known company will allow you to choose one of the alternatives shown in the following table . Your opportunity cost is 9%. a. Find the value today of each alternative. b. Are all the alternatives acceptable-that is, worth at least $23,000 today? c. Which alternative, if any, will you take? a. The present value of Alternative A is sRound to the nearest cent) The present value of Alternative B is (Round to the nearest cent) The present value of Alternative C is S b. Are all the alternatives acceptablethat is, worth at least $23,000 today? (Select the best answer below.) (Round to the nearest cent.) O Yes O No c. Which alternative, if any, will you take? (Select the best answer below.) O A. None OB. Alternative A O C. Alternative B O D. Alternative C
Answer #1

a)

Calculation pf present value of each alternative:

Alternative A

Opportunity cost = 9%

Cash inflow = $27500 at the end of Year 3

Present value of alternative A = $27500*Present value interest factor(9%,3) = $27500*0.77218 = $21234.95 i.e. $21235

Alternative B

Opportunity cost = 9%

Cash inflow = $53500 at the end of Year 9

Present value of alternative B = $53500*Present value interest factor(9%,9) = $53500*0.46042 = $24632.47 i.e. $24633

Alternative C

Opportunity cost = 9%

Cash inflow = $172000 at the end of Year 20

Present value of alternative C = $172000*Present value interest factor(9%,20) = $172000*0.17843 = $30689.96 i.e. $30690

b)

No, all the alternatives are not acceptable. Alternative A is not worth $23000 today.

c)

D. Alternative C

Alternative C is acceptable because it provides the highest present value.